OpenStudy (anonymous):

Divide and then reduce to lowest terms. 5/8 ÷ 3/4 please show me every step thx :)

OpenStudy (antonio_xx2):

5/8=0.625 and 3/4=0.75 and 0.625/0.75= 0.83333333333